# To appear in the 34th USENIX Security Symposium, August 13–15, 2025 DarkGram: A Large-Scale Analysis of Cybercriminal Activity Channels on Telegram

### v2

@@ 1 Introduction
    Finally, to further aid the research community in identifying malicious CACs, we have open-sourced our dataset— the first of its kind—along with the DarkGram framework, available at https://github.com/UTA-SPRLab/DarkGram.
@@ 3 Research Methodology
Harm to subscribers: Traditional cybercriminal forums, while serving as hubs for sharing malicious content, are also rife with bad actors who exploit their peers by sharing scams or harmful material that directly targets forum members [83, 36]. Section 5 investigates whether CAC administrators
exploit the trust of their audience by sharing content that can be harmful to the subscribers. We systematically use tools, such as VirusTotal and Hybrid Analysis, to identify phishing scams and malware disguised within the shared content.
@@ 4.2 Pirated software
To extend the software categorization across all the 11,800 posts identified by the classifier, we utilized the GPT-4 API [1] to automatically classify files based on the post text and the filename (when an attachment was present) within the existing 15 categories, with the potential to identify additional categories. We include the prompt used for this categorization in Section 10.1 in the Appendix. Using this approach, GPT-4 added four new categories, bringing the total number of categories to 19. Our coders validated the post category and description for a random sample of 1,000 such posts and found 994 to be accurate. Figure 3 illustrates the ten most common categories of software that appeared in these channels. Using GPT-4, we also identified that the 11,800 posts shared 6,154 applications, with 4,201 being for Android and 1,953 for Windows. In the majority of the posts (93.8%, 11,074), the software was shared directly as a file upload, though 2,144 of them also contained additional website links. Evaluating 100 of these samples manually, we found that these extra links were "backups" on an external website or cloud service. The remaining 726 posts did not have a file upload but relied on an external link (417), which also contained backup links. For example, a message might include multiple links, such as "Download link: [Shortened URL]. Backup link: [Another URL]." This redundancy ensures that users can access the software even if one link is taken down. Finally, for 309 posts, the users were directed to interact with one of 51 unique bots that had characteristics similar to those found previously Credential Compromise CACs, and discussed in Section 10.2 in the Appendix.
@@ 4.5 Social media manipulation
Methods of Building Trust and Advertising: Beyond enabling the purchase of fake engagement, 39 channels shared images as "proof" of successful services, likely to build trust with potential customers. These images showcased tangible results, such as increased engagement or follower counts, demonstrating the effectiveness of the services provided. For instance, Figure 6 highlights a channel displaying proof of a successful Instagram follower purchase, showing a visibly boosted follower count. Additionally, 45 channels frequently hosted time-limited giveaway competitions, offering their services for free to winners. Notably, three channels accounted for a significant number of giveaway posts (n=480) featuring collaborations with social media influencers boasting substantial followings. These posts often guaranteed a specific number of new followers for giveaway winners. For example, one post promoted a giveaway by an influencer with 9.1 million followers, promising winners between 10,000 and 15,000 new followers. These giveaways likely served multiple purposes: attracting new users through risk-free trials, increasing the visibility and engagement of the hosting channels, and generating urgency and excitement around the services. Collaborations with popular influencers further amplified the perceived legitimacy and appeal of the services, benefiting both the influencers and the service providers. Such strategies align with findings in related research, such as the promotion of NFTs [68], where scammers exploit artificial promotion to lure unsuspecting victims with promises of high returns or exclusive benefits. In Section 6, we explore how these "proofs" and giveaways contribute to enhanced engagement on these channels.
@@ 6 Channel Engagement
Impact of Post Forwards Post forwarding is a critical driver of channel engagement on Telegram, enabling content to be shared across channels and significantly expanding its reach. To evaluate the impact of forwards on subscriber growth, we analyzed forward activity in CAC posts and correlated it with subscriber growth over a seven-day period. Our findings reveal a strong positive association between the number of forwards and subscriber growth. Channels with posts that were forwarded more than 100 times in a week exhibited a median growth rate of 27.4%, compared to 12.1% for channels with fewer forwards ($p<0.01$). This trend was particularly pronounced in Pirated software and media channels. For instance, Pirated media channels with posts having over 100 forwards achieved a median growth rate of 39.2%, compared to 17.3% for channels with fewer forwards ($p<0.001$). Similarly, pirated software channels saw a median growth rate of 31.9% for posts forwarded over 100 times, compared to 14.2% for channels with lower forward activity ($p<0.001$). The nature of content shared in these channels plays a pivotal role in driving forward activity. Pirated media channels often feature exclusive or newly released materials that subscribers are eager to share, increasing the channel’s visibility and attracting new followers. Similarly, Pirated software channels frequently share links to high-demand software or updates, further amplifying their appeal and leading to substantial subscriber growth.
@@ 6.1 Emoji reactions
A majority of posts in the CAC channels (81.3%) were either closed to replies or did not receive any, and the replies that did exist were typically very brief, with a median length of only 4 words and an average of 9.16 words. This brevity stands in stark contrast to the more detailed replies [30] commonly seen on other social media platforms like Twitter (now X) and Facebook, making interpretation more challenging. Therefore, we instead focus on interpreting the emoji reactions to the posts in this section. Emojis provide a quick and expressive way for users to convey emotions and reactions without relying on lengthy text [67]. In CACs, analyzing emoji reactions offers valuable insights into user engagement and sentiment toward shared content. Emojis can reflect a range of emotions, such as approval, excitement, skepticism, or disapproval. By analyzing the frequency and types of emojis used in response to CAC posts, we can better understand the effectiveness, popularity, and overall community sentiment to the shared content.
@@ 7 Interpreting cross-platform dynamics
Accessibility to content: Cybercriminal forums are traditionally structured around multi-threaded, discussion-heavy ecosystems that rely on user-to-user interactions to foster trust, reputation, and mentoring [56, 49]. Accessing relevant material often requires browsing subforums, navigating multiple contributors, and becoming familiar with specialized jargon. Some forums also impose barriers—such as invite-only registrations or tiered “VIP” access levels—to maintain the exclusivity and quality of shared content [49, 63]. We also specifically noticed this in Section 4.7 where most blackhat resources on these forums where behind tiered barriers. While these mechanisms help sustain high standards among seasoned members, newcomers can find access to content daunting. In stark contrast, Telegram’s CACs operate on a broadcast model, wherein administrators share content via a simple chronological feed. Telegram CACs largely position subscribers as passive consumers, with minimal two-way engagement beyond emoji reactions or brief replies. Additionally, each Telegram CAC typically focuses on a single type of content, streamlining the user experience by eliminating the need for extensive browsing or vetting.
Heightened Subscriber Risk: We argue that subscribers face greater risks on Telegram CACs’ channels than traditional cybercrime forums. Research on such forums reveals a surprising degree of self-regulation [27], where experienced members act as informal gatekeepers by spot-checking questionable posts and leaving negative feedback to deter low-quality or harmful contributions [56] - mechanisms which help to mitigate risks for members. In contrast, our analysis of Telegram CACs an environment which can be dangerous to the subscribers. Specifically, in Section 5, we find 28.1% of URLs shared on these channels led to phishing sites, and 38% of shared executable files were malicious. Unlike forums, where scams/harmful content typically unfold over days or weeks, allowing members to warn others [52], administrators on Telegram CACs can inject such content into the feed almost instantly, and since the majority of the CACs disable replies to posts, wary subscribers have no opportunity to raising alarms then they notice suspicious activity. It was also alarming that a significant portion of such harmful material is shared in channels focused on pirated software, media, and social media manipulation - content that is seeked by regular users [12] instead of technically proficient hackers . This is further evidenced by our findings in Section 6 where users often react positively to posts with malicious content, suggesting a lack of awareness of the risks in accessing them.
@@
Takeaway: Prior literature has identified how compromised credentials [16], SMM panels [58] and blackhat resources [70] are regularly shared on cybercriminal forums. Our analysis in Section 4.7 further reinforces these findings, revealing that forums not only exhibit similar payload characteristics but also rely on overlapping external websites to distribute their content. This highlights that Telegram CACs, at the very least, function as an alternative source for accessing cybercriminal content. Additionally, Telegram CACs share pirated content, which is largely absent from traditional cybercriminal forums. What sets Telegram apart is its ability to deliver this content in a streamlined manner that is highly resilient to takedown efforts and readily accessible to a broad audience. Our findings also reveal that this content not only reaches and is actively consumed by millions of users, not only necessitating the need for their swift takedown but also as a resource for security vendors for identifying malicious content. In the next section, we demonstrate how DarkGram, our BERT-based detection pipeline, can be extended to identify and disrupt newly emerging CACs in real-time.
@@ 8 Detecting New CACs Using DarkGram
We extended the DarkGram framework to proactively identify new CACs shared on Telegram and Facebook from May 26 to August 11, 2024. Our goal was to detect these channels early, report them to Telegram for removal before they could gain significant subscribers, and notify affected stakeholders and security vendors regarding the malicious content shared within these channels. Similar to how we reported the original CACs in Section 4, DarkGram also automatically reports to Telegram’s abuse email with the channel description, CAC category, and screenshots of the post. To identify new CACs on Telegram, we first searched for URLs starting with t.me shared across the 339 CACs, either as a post or through a bot. This approach was chosen because the channels in our ground truth already represent a curated, high-risk dataset, making it both resource-efficient and effective to focus on the interconnected networks commonly observed in cybercriminal activity [45]. As described in Section 6, "Channel Migration," CACs often promote and encourage users to join similar channels. By leveraging this observed behavior, our approach can increase the speed and likelihood of detecting new emerging CACs. Nevertheless, DarkGram can just as easily be run on a random sample of channels with equal ease. For each t.me link, DarkGram analyzes the 10 most recent posts to emulate the methodology we used to identify malicious CACs from Telemetrio while building the ground truth (Section 2.1). Depending on the flagged content, DarkGram extracted URLs and executable files for further analysis using VirusTotal [11] and PhishIntention. Channels with five or more flagged posts were classified as malicious and reported. Building the ground-truth in Section 2.1 also informed this threshold, as legitimate channels would occasionally contain malicious posts, likely shared accidentally by the administrators, a phenomenon not uncommon in other social media platform [14]. By setting a higher threshold, we ensure that only channels demonstrating consistent malicious behavior are flagged, reducing the risk of false positives. Using this approach, DarkGram identified 127 malicious channels out of 245 examined links. To demonstrate that DarkGram can also be used to find CACs on other social media platforms, we run it on Facebook, which is often used by cybercriminals to share malicious content [20]. To identify CACs on Facebook, we utilized the CrowdTangle API [3] (now discontinued and replaced by the Meta Content Library). Through CrowdTangle’s global search functionality, DarkGram extracted t.me links embedded in posts from public groups. These links were evaluated using the same process as links found on Telegram CACs, looking through 4,191 links from 3,002 groups and flagged 69 (shared in 14 groups) as malicious, which were reported.
@@
Characteristics: Overall, the 196 channels shared on Telegram and Facebook combined were spread across all CAC categories we have studied: Credential Compromise (56), Pirated software (72), Artificial Boosting (17), Pirated Media (23), and Blackhat Resources (28). Figure 8 in Appendix illustrates the distribution of the new channels found by DarkGram over the three months. Manually evaluating a maximum of 10 randomly selected posts from each of the 196 newly discovered CACs, we found that the content shared across them closely mirrors the patterns observed in the 339 ground-truth CAC channels. These similarities span all CAC categories, including tactics employed by channel owners such as using "proofs" and bots for content distribution. For brevity, we do not go into the detailed statistics of this content. However, this finding underscores the persistence of the malicious activities that we have previously documented in this paper extending to the newer channels as well. Based on the creation dates of the channels, we also found that the new channels had a median channel age of only 27 days compared to the median age of our original CAC channels at 237 days. This indicates that DarkGram was able to find significantly newer channels, as also justified by the newer channels having a median follower count of only 1,408.
@@
Disclosure and Removal: Reporting these channels to Telegram led to the removal of all 196 channels, with a median response time of 4 days. This is in stark contrast to when we reported the original channels in Section 4.6 where only 64 (out of 339) channels were removed. With Telegram’s content removal policy not transparent on how reported channels are moderated, we assume that since the original CACs were more established with higher subscriber counts and more traffic, Telegram might be more resistant towards removing them, as is also the practice on other social media platforms where scammers taking over popular accounts are more resilient than newer accounts [24]. This further brings into focus the importance of identifying and removing the CACs early. We also reported 20 of the 23 false negatives missed by DarkGram (19 from Telegram links and 4 from Facebook) at the end of the study. By that time, these channels had a median subscriber count of 8,125, significantly higher than when they were initially scanned by DarkGram. The three channels that were not reported, was since they had already been removed/takedown. Our disclosure resulted in the removal of 14 channels (as communicated by Telegram and verified by us), while 6 remained active. It thus appears that Telegram’s urgency in removing the channels becomes less severe the more popular the channel becomes, further emphasizing the need for early detection, and DarkGram appears to have a positive impact in identifying and removing these channels. This is in line with recent reports of Telegram’s struggles to stop cybercriminal activity from running rampant on their service [44, 72], while the reports focus on cybercriminal activity in private chats, through this work we see that malicious activity is just as functional on public easily accessible channels and often without any resistance from Telegram. With respect to malicious content found on the new CACs, we reported 597 malicious URLs and 314 files, resulting in the removal of 343 URLs, 157 of which were confirmed by domain providers as a direct result of our reports. The DarkGram framework is open-sourced at https://github.com/UTA-SPRLab/DarkGram.

### v3

# DarkGram: A Large-Scale Analysis of Cybercriminal Activity Channels on Telegram

@@ 1 Introduction
    Finally, to further aid the research community in identifying malicious CACs, we have open-sourced our dataset, the first of its kind, along with the DarkGram framework, available at https://github.com/Scalable-Security-Research-Lab/DarkGram.
@@ 3 Research Methodology
Harm to subscribers: Traditional cybercriminal forums, while serving as hubs for sharing malicious content, are also rife with bad actors who exploit their peers by sharing scams or harmful material that directly targets forum members [88, 40]. Section 5 investigates whether CAC administrators exploit the trust of their audience by sharing content that can be harmful to the subscribers. We systematically use tools, such as VirusTotal and Hybrid Analysis, to identify phishing scams and malware disguised within the shared content.
@@ 4.2 Pirated software
To extend the software categorization across all the 11,800 posts identified by the classifier, we utilized the GPT-4 API [1] to automatically classify files based on the post text and the filename (when an attachment was present) within the existing 15 categories, with the potential to identify additional categories. We include the prompt used for this categorization in Section 11.1 in the Appendix. Using this approach, GPT-4 added four new categories, bringing the total number of categories to 19. Our coders validated the post category and description for a random sample of 1,000 such posts and found 994 to be accurate. Figure 3 illustrates the ten most common categories of software that appeared in these channels. Using GPT-4, we also identified that the 11,800 posts shared 6,154 applications, with 4,201 being for Android and 1,953 for Windows. In the majority of the posts (93.8%, 11,074), the software was shared directly as a file upload, though 2,144 of them also contained additional website links. Evaluating 100 of these samples manually, we found that these extra links were "backups" on an external website or cloud service. The remaining 726 posts did not have a file upload but relied on an external link (417), which also contained backup links. For example, a message might include multiple links, such as "Download link: [Shortened URL]. Backup link: [Another URL]." This redundancy ensures that users can access the software even if one link is taken down. Finally, for 309 posts, the users were directed to interact with one of 51 unique bots that had characteristics similar to those found previously in Credential Compromise CACs, and discussed in Section 11.2 in the Appendix.
@@ 4.5 Social media manipulation
Methods of Building Trust and Advertising: Beyond enabling the purchase of fake engagement, 39 channels shared images as "proof" of successful services, likely to build trust with potential customers. These images showcased tangible results, such as increased engagement or follower counts, demonstrating the effectiveness of the services provided. For instance, Figure 6 highlights a channel displaying proof of a successful Instagram follower purchase, showing a visibly boosted follower count. Additionally, 45 channels frequently hosted time-limited giveaway competitions, offering their services for free to winners. Notably, three channels accounted for a significant number of giveaway posts (n=480) featuring collaborations with social media influencers boasting substantial followings. These posts often guaranteed a specific number of new followers for giveaway winners. For example, one post promoted a giveaway by an influencer with 9.1 million followers, promising winners between 10,000 and 15,000 new followers. These giveaways likely served multiple purposes: attracting new users through risk-free trials, increasing the visibility and engagement of the hosting channels, and generating urgency and excitement around the services. Collaborations with popular influencers further amplified the perceived legitimacy and appeal of the services, benefiting both the influencers and the service providers. Such strategies align with findings in related research, such as the promotion of NFTs [71], where scammers exploit artificial promotion to lure unsuspecting victims with promises of high returns or benefits. In Section 6, we see how such "proofs" and giveaways contribute to enhanced engagement on these channels.
@@ 4.7 Overlap of content between Telegram CACs and cybercriminal forums
Annotation using Generative AI: We used GPT-4 for annotating data from posts shared in "Credential Compromise" and "Social Media Manipulation" CACs, depending on its proven effectiveness in similar tasks as highlighted in prior work [33, 81]. To ensure accuracy, we also manually evaluated a sample of the annotations in each instance. However, we acknowledge the reliance on a proprietary platform raises concerns about long-term replicability. In the event GPT-4 becomes unavailable or unsupported, the annotations it generated can be reproduced using open-source alternatives like LLaMA 3.1, which not only demonstrate comparable or superior performance to GPT-4 [10] but also benefit from a more recent knowledge cut-off dates.
@@ 6 Channel Engagement
Impact of Post Forwards Post forwarding is a critical driver of channel engagement on Telegram, enabling content to be shared across channels and significantly expanding its reach. To evaluate the impact of forwards on subscriber growth, we analyzed forward activity in CAC posts and correlated it with subscriber growth over a seven-day period. Our findings reveal a strong positive association between the number of forwards and subscriber growth. Channels with posts that were forwarded more than 100 times in a week exhibited a median growth rate of 27.4%, compared to 12.1% for channels with fewer forwards ($p<0.01$). This trend was particularly pronounced in Pirated software and media channels. For instance, Pirated media channels with posts having over 100 forwards achieved a median growth rate of 39.2%, compared to 17.3% for channels with fewer forwards ($p<0.001$). Similarly, pirated software channels saw a median growth rate of 31.9% for posts forwarded over 100 times, compared to 14.2% for channels with lower forward activity ($p<0.001$). The nature of content shared in these channels plays a pivotal role in driving forward activity. Pirated media channels often feature newly released materials that subscribers are eager to share, increasing the channel’s visibility and attracting new followers. Similarly, Pirated software channels share links to high-demand software or updates, further amplifying their appeal and leading to substantial subscriber growth.
@@ 6.1 Emoji reactions
A majority of posts in the CAC channels (81.3%) were either closed to replies or did not receive any, and the replies that did exist were typically very brief, with a median length of only 4 words and an average of 9.16 words. This brevity stands in stark contrast to the more detailed replies [34] commonly seen on other social media platforms like Twitter (now X) and Facebook, making interpretation more challenging. Therefore, we instead focus on interpreting the emoji reactions to the posts in this section. Emojis provide a quick and expressive way for users to convey emotions and reactions without relying on lengthy text [70]. In CACs, analyzing emoji reactions offers valuable insights into user engagement and sentiment toward shared content. Emojis can reflect a range of emotions, such as approval, excitement or disapproval. By analyzing the frequency and types of emojis used in response to CAC posts, we can better understand the effectiveness, popularity, and overall community sentiment to the shared content.
@@ 7 Interpreting cross-platform dynamics
Accessibility to content: Many cybercriminal forums are structured around multi-threaded, discussion-heavy ecosystems that rely on user-to-user interactions to foster trust, reputation, and mentoring [59, 52]. Accessing relevant material often requires browsing subforums, navigating multiple contributors, and becoming familiar with specialized jargon. Some forums also impose barriers—such as invite-only registrations or tiered “VIP” access levels—to maintain the exclusivity and quality of shared content [52, 66]. We also specifically noticed this in Section 4.7 where most blackhat resources on these forums where behind tiered barriers. While these mechanisms help sustain high standards among seasoned members, newcomers can find access to content daunting. In stark contrast, Telegram’s CACs operate on a broadcast model, wherein administrators share content via a simple chronological feed. Telegram’s CACs often present content in a format that requires minimal active engagement from subscribers. Additionally, each Telegram CAC typically focuses on a single type of content, streamlining the user experience by eliminating the need for extensive browsing or vetting.
Heightened Subscriber Risk: We argue that subscribers face potentially greater risks on Telegram CACs’ channels than traditional cybercrime forums. Research on such forums reveals a surprising degree of self-regulation [30], where experienced members act as informal gatekeepers by spot-checking questionable posts and leaving negative feedback to deter low-quality or harmful contributions [59] - mechanisms which help to mitigate risks for members. In contrast, our analysis of Telegram CACs an environment which can be dangerous to the subscribers. Specifically, in Section 5, we find 28.1% of URLs shared on these channels led to phishing sites, and 38% of shared executable files were malicious. Unlike forums, where scams/harmful content typically unfold over days or weeks, allowing members to warn others [55], administrators on Telegram CACs can inject such content into the feed almost instantly, and since the majority of the CACs disable replies to posts, wary subscribers have limited opportunity to raising alarms then they notice suspicious activity. It was also alarming that a significant portion of such harmful material is shared in channels focused on pirated software, media, and social media manipulation - content that is seeked by regular users [14] instead of technically proficient hackers . This is further evidenced by our findings in Section 6 where users often react positively to posts with malicious content, suggesting a lack of awareness of the risks in accessing them.
@@
Takeaway: Prior literature has identified how compromised credentials [19], SMM panels [61] and blackhat resources [73] are regularly shared on cybercriminal forums. Our analysis in Section 4.7 further reinforces these findings, revealing that forums not only exhibit similar payload characteristics but also rely on overlapping external websites to distribute their content. This highlights that Telegram CACs, at the very least, function as an alternative source for accessing cybercriminal content. Additionally, Telegram CACs share pirated content, which is largely absent from traditional cybercriminal forums. What sets Telegram apart is its ability to deliver this content in a streamlined manner that is readily accessible to a broad audience, and an apparent lack of strong moderation on the platform makes removal of such content less likely. Our findings also reveal that this content not only reaches and is actively consumed by millions of users, not only necessitating the need for their swift takedown but also as a resource for security vendors for identifying malicious content. In the next section, we demonstrate how DarkGram, our BERT-based detection pipeline, can be extended to identify and disrupt newly emerging CACs in real-time.
@@ 8 Detecting New CACs Using DarkGram
We extended the DarkGram framework to proactively identify new CACs shared on Telegram and Facebook from May 26 to August 11, 2024. Our goal was to detect these channels early, report them to Telegram for removal before they could gain significant subscribers, and notify affected stakeholders and security vendors regarding the malicious content shared within these channels. Similar to how we reported the original CACs in Section 4, DarkGram also automatically reports to Telegram’s abuse email with the channel description, CAC category, and screenshots of the post. To identify new CACs on Telegram, we first searched for URLs starting with t.me shared across the 339 CACs, either as a post or through a bot. This approach was chosen because the channels in our ground truth already represent a curated, high-risk dataset, making it both resource-efficient and effective to focus on the interconnected networks commonly observed in cybercriminal activity [49]. As described in Section 6, "Channel Migration," CACs often promote and encourage users to join similar channels. By leveraging this observed behavior, our approach can increase the speed and likelihood of detecting new emerging CACs. Nevertheless, DarkGram can just as easily be run on a random sample of channels with equal ease. For each t.me link, DarkGram analyzes the 10 most recent posts to emulate the methodology we used to identify malicious CACs from Telemetrio while building the ground truth (Section 2.1). Depending on the flagged content, DarkGram extracted URLs and executable files for further analysis using VirusTotal [13] and PhishIntention. Channels with five or more flagged posts were classified as malicious and reported. Building the ground-truth in Section 2.1 also informed this threshold, as legitimate channels would occasionally contain malicious posts, likely shared accidentally by the administrators, a phenomenon not uncommon in other social media platform [17]. By setting a higher threshold, we ensure that only channels demonstrating consistent malicious behavior are flagged, reducing the risk of false positives. Using this approach, DarkGram identified 127 malicious channels out of 245 examined links. To demonstrate that DarkGram can also be used to find CACs on other social media platforms, we run it on Facebook, which is often used by cybercriminals to share malicious content [23]. To identify CACs on Facebook, we utilized the CrowdTangle API [3] (now replaced by the Meta Content Library). Through CrowdTangle’s global search functionality, DarkGram extracted t.me links embedded in posts from public groups. These links were evaluated using the same process as links found on Telegram CACs, looking through 4,191 links from 3,002 groups and flagged 69 (shared in 14 groups) as malicious, which were reported.
@@
Characteristics: Overall, the 196 channels shared on Telegram and Facebook combined were spread across all CAC categories we have studied: Credential Compromise (56), Pirated software (72), Artificial Boosting (17), Pirated Media (23), and Blackhat Resources (28). Figure 8 in Appendix illustrates the distribution of the new channels found by DarkGram over the three months. Manually evaluating a maximum of 10 randomly selected posts from each of the 196 newly discovered CACs, we found that the content shared across them closely mirrors the patterns observed in the 339 ground-truth CAC channels. These similarities span all CAC categories, including tactics employed by channel owners, such as using "proofs" and bots for content distribution. For brevity, we do not go into the detailed statistics of this content. However, this finding underscores the persistence of the malicious activities previously documented in this paper extending to the newer channels as well. Based on the creation dates of the channels, we found that the new channels had a median channel age of only 27 days compared to that of our original CAC channels at 237 days. This indicates that DarkGram was able to find significantly newer channels, as also justified by the newer channels having a median follower count of only 1,408.
@@
Disclosure and Removal: Reporting these channels to Telegram led to the removal of all 196 channels, with a median response time of 4 days. This is in contrast to when we reported the original channels in Section 4.6 where only 64 (of 339) channels were removed. With Telegram’s content removal policy not transparent on how reported channels are moderated, we assume that since the original CACs were more established with higher subscriber counts and more traffic, Telegram might be more resistant towards removing them, as is also the practice on other social media platforms where scammers taking over popular accounts are more resilient than newer accounts [27]. This further brings into focus the importance of identifying and removing the CACs early. We also reported 20 of the 23 false negatives missed by DarkGram (19 from Telegram links and 4 from Facebook) at the end of the study. By that time, these channels had a median subscriber count of 8,125, significantly higher than when they were initially scanned by DarkGram. The three channels that were not reported, was since they had already been removed/takedown. Our disclosure resulted in the removal of 14 channels (as communicated by Telegram and verified by us), while 6 remained active. It thus appears that Telegram’s urgency in removing the channels becomes less severe the more popular the channel becomes, further emphasizing the need for early detection, and DarkGram appears to have a positive impact in identifying and removing these channels. This is in line with recent reports of Telegram’s struggles to stop cybercriminal activity from running rampant on their service [48, 75], while the reports focus on cybercriminal activity in private chats, through this work we see that malicious activity is just as functional on public easily accessible channels and often without any resistance from Telegram. We also reported 597 malicious URLs and 314 files, resulting in the removal of 343 URLs, 157 of which were confirmed by domain providers as a direct result of our reports.